10. If the moment-generating function of X is find the mean. variance. and omf of X.
Answer #1

as we know that moment generating function Mx(t) =\sumP(x)etx

comparing given mgf to above:

\below is pmf of X:

P(X=1)=2/5

P(X=2)=1/5

P(X=3)=2/5

mean E(X) =\sumxP(x)=1*(2/5)+2*(1/5)+3*(2/5)=10/5=2

E(X2)=\sumx2P(x)=12*(2/5)+22*(1/5)+32*(2/5)=24/5=4.8

hence Variance =E(X2)-(E(X))2 =4.8-22 =0.8